# Beyond Joint Demonstrations: Personalized Expert Guidance for Efficient Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Multi-Agent Reinforcement Learning (MARL) algorithms face the challenge of efficient exploration due to the exponential increase in the size of the joint state-action space. While demonstration-guided learning has proven beneficial in single-agent settings, its direct applicability to MARL is hindered by the practical difficulty of obtaining joint expert demonstrations. In this work, we introduce a novel concept of personalized expert demonstrations, tailored for each individual agent or, more broadly, each individual type of agent within a heterogeneous team. These demonstrations solely pertain to single-agent behaviors and how each agent can achieve personal goals without encompassing any cooperative elements, thus naively imitating them will not achieve cooperation due to potential conflicts. To this end, we propose an approach that selectively utilizes personalized expert demonstrations as guidance and allows agents to learn to cooperate, namely personalized expert-guided MARL (PegMARL). This algorithm utilizes two discriminators: the first provides incentives based on the alignment of policy behavior with demonstrations, and the second regulates incentives based on whether the behavior leads to the desired objective. We evaluate PegMARL using personalized demonstrations in both discrete and continuous environments. The results demonstrate that PegMARL learns near-optimal policies even when provided with suboptimal demonstrations, and outperforms state-of-the-art MARL algorithms in solving coordinated tasks. We also showcase PegMARL’s capability to leverage joint demonstrations in the StarCraft scenario and converge effectively even with demonstrations from non-co-trained policies.
@@ 1 Introduction
The use of expert demonstrations has been proven effective in accelerating learning in single-agent reinforcement learning, as evidenced by studies such as Kang et al. (2018); Chen & Xu (2022); Rengarajan et al. (2022). This approach has since been extended to Multi-Agent Reinforcement Learning (MARL) (Lee & Lee, 2019; Qiu et al., 2022), which typically assumes the availability of high-quality collaborative joint demonstrations. However, from a practical standpoint, collecting joint demonstrations can be labor-intensive, demanding one user per agent in cooperative scenarios. Furthermore, these demonstrations are not scalable. If we change the number of agents or introduce new types of agents, we will need to gather a new set of demonstrations to learn from.
In contrast, it is much easier to obtain demonstrations for individual agents, or even better, for each type of agents in a heterogeneous setting. We thus ask the following research question: could we leverage individual-wise task demonstrations instead? In this work, we refer to such expert demonstrations that address single-agent behaviors for personal objectives as personalized demonstrations (Figure 1). Since the personalized demonstrations will not necessarily reflect how the agents can collaborate and may even conflict with each other in the joint setting, naively mimicking the demonstrations will not achieve cooperation. Therefore, purely imitation learning-based approaches would not be effective. We need an approach that selectively utilizes suitable personalized expert demonstrations as guidance and allows agents to learn to cooperate via collecting reward signals from the environments.
@@
To this end, we present our algorithm, Personalized Expert-Guided MARL (PegMARL), which carries out reward-shaping as a form of guidance. We implement this via two discriminators. The first, a personalized behavior discriminator, evaluates local state-action pairs, providing positive incentives for actions that align with the demonstration and negative incentives for divergent ones. The second, a personalized dynamics discriminator, assesses whether a local state-action pair induces a desired change in dynamics similar to that observed in the demonstration, adjusting the incentive weight accordingly. We prove the effectiveness of PegMARL on both discrete gridworld and continuous multi-agent particle environments (Lowe et al., 2017; Mordatch & Abbeel, 2017). The main contributions of this paper are as follows:
(1) Our approach is the first to enable utilizing personalized demonstrations for policy learning in heterogeneous MARL environments, regardless of the number and type of agents involved.
(2) We propose PegMARL that dynamically and selectively reshapes the original reward to aid exploration. It is worth mentioning that this algorithm is general and compatible with most MARL policy gradient methods.
(3) We evaluate PegMARL with personalized demonstrations in both discrete and continuous environments. Our algorithm outperforms state-of-the-art decentralized MARL algorithms, pure multi-agent imitation learning, and reward-shaping techniques in terms of scalability and convergence speed, and obtains near-optimal policies even when provided with suboptimal demonstrations.
(4) We showcase PegMARL’s capability to also leverage joint demonstrations, regardless of whether they are sampled from co-trained or non-co-trained policies111“Co-trained policies” refers to policies that have been trained together in the same environment with shared experiences. We refer the readers to Section 6.2 of Wang et al. (2023) for more information.. Experimental results on the StarCraft environment (Samvelyan et al. (2019)) demonstrate that PegMARL converges effectively even with demonstrations from non-co-trained policies.
@@ 2 Related Works
Imitation Learning (IL). IL methods seek to replicate an expert policy from demonstration data generated by that policy. Behavior Cloning (BC) is a commonly employed IL technique (Pomerleau (1991); Bojarski et al. (2016)), where the expert policy is estimated through supervised learning on demonstration data. However, BC is susceptible to compounding errors resulting from distribution shifts (Ross et al. (2011)). Another thread of IL research is Inverse Reinforcement Learning (IRL) (Ng et al. (2000); Ziebart et al. (2008)), in which the underlying reward function is estimated from the demonstration data and then used for policy learning. To alleviate the computational overhead of IRL, Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon (2016)) was introduced, allowing direct policy learning from observed data without intermediate IRL steps. Song et al. (2018) extended this approach to introduce Multi-Agent GAIL (MAGAIL), which adapts GAIL to high-dimensional environments featuring multiple cooperative or competing agents. In general, IL approaches can rarely perform better than demonstrations. Therefore, they are not directly suitable for scenarios where only personalized expert demonstrations that do not demonstrate how to collaborate are available.
@@
Learning from Demonstration (LfD). In contrast to IL methods, LfD aims to leverage demonstration data to facilitate learning rather than simply mimicking expert behavior. Existing LfD works incorporate demonstration data into a replay buffer with a prioritized replay mechanism to accelerate the learning process. Due to the off-policy nature of the demonstration data, most methods are value-based Hester et al. (2018); Vecerik et al. (2017). There has been some recent work where the demonstrations are used to aid exploration, especially in environments with large state-action spaces (Kang et al. (2018); Chen & Xu (2022); Rengarajan et al. (2022)). For instance, POfD (Kang et al. (2018)) learns an implicit reward from the demonstration data using a discriminator and incorporates it into the original sparse reward. LOGO (Rengarajan et al. (2022)) uses the demonstration data to directly guide the policy update: during each update iteration, the algorithm seeks a policy that closely resembles the behavior policy within a trust region. However, these approaches are primarily focused on single-agent settings.
@@
In multi-agent settings, Qiu et al. (2022) suggests using demonstrations to pretrain agents through imitation learning as a warm start, followed by optimization of the pretrained policies using standard MARL algorithms. Lee & Lee (2019) augment the experience buffer with demonstration trajectories and gradually decrease the mixing of demonstration samples during training to prevent the learned policy from being overly influenced by demonstrations. Similar to POfD, DM2 (Wang et al. (2023)) enables agents to enhance their task-specific rewards by training discriminators as well. Each agent matches toward a target distribution of concurrently sampled trajectories from a joint expert policy to facilitate coordination. These approaches, however, require joint demonstrations of the same team configurations, which can be cumbersome (as we mentioned in section 1). Our approach differs in that we leverage only personalized expert demonstrations to learn a cooperative policy.
@@ 3 Preliminaries
We start by considering a Markov Decision Process $(\mathcal{S},\mathcal{A},P,r,\gamma)$ for a cooperative multi-agent setting with $N$ agents. Here, $\mathcal{S}$ denotes the global state across all the agents, which can be decomposed as the product of $N$ local spaces $\mathcal{S}_{i}$ as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}\times\cdots\times\mathcal{S}_{N}$. By noting the local state of agent $i$ as $s_{i}\in\mathcal{S}_{i}$, the global state is $s=(s_{1},s_{2},\cdots,s_{N})$. Similarly, we define the global action space $\mathcal{A}$ as $\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{N}$, meaning that for any $a\in\mathcal{A}$, we may write $a=(a_{1},a_{2},\cdots,a_{N})$ with $a_{i}\in\mathcal{A}_{i}$. The transition probability from state $s$ to $s^{\prime}$ after taking a joint action $a$ is denoted by $P(s^{\prime}|s,a)=\prod_{i}P_{i}(s_{i}^{\prime}|s,a)$. All agents share a common reward function $r$, and $\gamma\in[0,1]$ is the discount factor. In this work, we focus on decentralized learning and define the global policy as $\pi_{\theta}(a|s)$, where $\theta\in\Theta$ are the policy parameters. Specifically, we have $\theta=(\theta_{1},\theta_{2},\cdots,\theta_{N})$ as the factorized global policy parameters, and we can write $\pi_{\theta}(a|s)=\prod_{i}\pi_{\theta_{i}}(a_{i}|s)$ using policy factorization.
@@
Personalized Tasks and MDPs. To introduce the notion of personalized demonstration, we need to extract the individual tasks of each agent from the collective tasks of multiple agents. For example, in Figure 2, the green agent’s personalized task is to open the door and the red agent’s personalized task is to reach the key without the other’s presence. We then define Personalized Markov Decision Processes (PerMDPs) $(\mathcal{S}_{i},\mathcal{A}_{i},Q_{i},r_{i},\gamma)$ for each agent or, more generally, for each $type$ of agents that share the same objective. Here, a $type$ refers to behaviorally identical agents using the formalism from Bettini et al. (2023), and we provide the PerMDP definition for each agent for simplicity. We assume that the state space $\mathcal{S}_{i}$ and action space $\mathcal{A}_{i}$ of the personalized task $\mathcal{T}_{i}$ are the same as the local state and action spaces of agent $i$ from the joint task. The transition probability from state $s_{i}$ to $s_{i}^{\prime}$ after taking a action $a_{i}$ is represented as $Q_{i}(s^{\prime}_{i}|s_{i},a_{i})$. By following an arbitrary policy $\pi_{i}(a_{i}|s_{i})$ for this personalized task, we can collect a set of personalized demonstrations $\mathcal{B}_{i}=\{(s_{i}^{t},a_{i}^{t})\}_{t=0}^{H}$, where $H$ is the episode horizon. We’d like to emphasize that while we assume the personalized tasks and the joint task are conducted in the same environment map, the underlying transition dynamics are different. Additionally, the joint reward $r$ may not necessarily equal the summation of rewards $r_{i}$ for each personalized task.
@@ 4 MARL with Personalized Expert Demonstrations
Now, we are ready to present the main problem we are interested in solving in this work. Assume each agent $i$ is associated with a personalized task $\mathcal{T}_{i}$. We collect one set of expert demonstrations for each agent $i$ or, equivalently, each personalized task $\mathcal{T}_{i}$. By letting an expert user perform each personalized task in the respective personalized MDP, we obtain a collection of expert demonstrations denoted by $\{\mathcal{B}_{E_{1}},\mathcal{B}_{E_{2}},...,\mathcal{B}_{E_{N}}\}$. We assume that the underlying expert policy associated with $\mathcal{B}_{E_{i}}$ is $\pi_{E_{i}}(a_{i}|s_{i})$, and $\lambda^{\pi_{E_{i}}}$ is the occupancy measure following the expert’s policy $\pi_{E_{i}}$ for agent $i$.
@@ 4.1 Formulating Personalized Expert-Guided MARL
In standard Multi-Agent Reinforcement Learning, agents aim to discover optimal joint policies that maximize the long-term return $R(\pi_{\theta})\mathrel{\mathop{\ordinarycolon}}=\langle\lambda^{\pi_{\theta}},r\rangle=\frac{1}{N}\sum_{i=1}^{N}\langle\lambda^{\pi_{\theta}}_{i},r\rangle$. Typically, the learning process commences with random exploration, which is often inefficient due to MARL’s exponentially growing exploration spaces, especially when rewards are sparse. Beginning with the intuition of leveraging personalized demonstrations as guidance for how each agent should accomplish their personalized tasks to promote more effective exploration, akin to Kang et al. (2018) for single-agent cases, we can define the objective for learning from personalized demonstrations in multi-agent settings as follows:
@@
However, blindly imitating the personalized demonstration may not always yield favorable outcomes and can even impede the learning process. Previous works in MARL have predominantly utilized joint demonstrations as guidance. As demonstrated by DM2 (Wang et al. (2023)), trajectories must be sampled from a co-trained joint expert policy for their joint action-matching objective to converge. The crucial aspect that joint demonstrations offer, which personalized demonstrations lack, is compatibility. To illustrate, consider the example depicted in Figure 3, where two agents aim to swap positions without entering the danger region. Two sets of optimal joint policies exist for this task: agent $a$ takes path 1 and agent $b$ takes path 2, or vice versa. When sampling joint demonstrations from a set of co-trained policies, the agents’ behaviors will be naturally compatible. In contrast, if personalized demonstrations were provided, agents could indifferently choose between both paths to reach their goals, potentially resulting in conflicting strategies through naive imitation.
@@
Consider the scenario depicted in Figure 4, where the blue agent seeks to replicate its personalized demonstration within a multi-agent environment. The blue agent can successfully transition to the neighboring grid on the right if it’s unoccupied; otherwise, the transition fails. Viewing the blue agent as the focal agent and the green agent as part of the environment, successful imitation hinges on the agent’s “local transition dynamics” matching those of the environment where the personalized demonstrations originated. We define $P(s_{i}^{\prime}|s_{i},a_{i})$ as an approximation for the $i$-th agent’s local transition dynamics in the multi-agent environment, which is governed by the true transition dynamics $P(s^{\prime}|s,a)$ and the policy $\pi_{\theta}$. With this insight, we further refine the objective function as follows:
@@
where $Q(s_{i}^{\prime}|s_{i},a_{i})$ represents the true transition dynamics of the personalized MDP from which demonstrations are collected. By adjusting the learning objective in this manner, the occupancy matching only happens on those state-action pairs that guide us toward the desired next local state under the current policy.

### v3

@@ Abstract
Multi-Agent Reinforcement Learning (MARL) algorithms face the challenge of efficient exploration due to the exponential increase in the size of the joint state-action space. While demonstration-guided learning has proven beneficial in single-agent settings, its direct applicability to MARL is hindered by the practical difficulty of obtaining joint expert demonstrations. In this work, we introduce a novel concept of personalized expert demonstrations, tailored for each individual agent or, more broadly, each individual type of agent within a heterogeneous team. These demonstrations solely pertain to single-agent behaviors and how each agent can achieve personal goals without encompassing any cooperative elements, thus naively imitating them will not achieve cooperation due to potential conflicts. To this end, we propose an approach that selectively utilizes personalized expert demonstrations as guidance and allows agents to learn to cooperate, namely personalized expert-guided MARL (PegMARL). This algorithm utilizes two discriminators: the first provides incentives based on the alignment of individual agent behavior with demonstrations, and the second regulates incentives based on whether the behaviors lead to the desired outcome. We evaluate PegMARL using personalized demonstrations in both discrete and continuous environments. The experimental results demonstrate that PegMARL outperforms state-of-the-art MARL algorithms in solving coordinated tasks, achieving strong performance even when provided with suboptimal personalized demonstrations. We also showcase PegMARL’s capability of leveraging joint demonstrations in the StarCraft scenario and converging effectively even with demonstrations from non-co-trained policies.
@@ 1 Introduction
The use of expert demonstrations [^1] In our context, the term “expert guidance” refers to demonstrations that provide meaningful, above-random performance in complex environments, even if the demonstrations are not necessarily optimal. has been proven effective in accelerating learning in single-agent reinforcement learning, as evidenced by studies such as Kang et al. (2018); Chen & Xu (2022); Rengarajan et al. (2022). This approach has since been extended to Multi-Agent Reinforcement Learning (MARL) (Lee & Lee, 2019; Qiu et al., 2022), which typically assumes the availability of high-quality collaborative joint demonstrations. However, from a practical standpoint, collecting joint demonstrations can be labor-intensive, demanding one user per agent in cooperative scenarios. Furthermore, these demonstrations are not scalable. If we change the number of agents or introduce new types of agents, we will need to gather a new set of demonstrations to learn from.
In contrast, it is much easier to obtain demonstrations for individual agents, or even better, for each type of agents in a heterogeneous setting. We thus ask the following research question: could we leverage individual-wise task demonstrations instead? In this work, we refer to such expert demonstrations that address single-agent behaviors for personal objectives as personalized demonstrations (see Figure 1). Since the personalized demonstrations will not necessarily reflect how the agents can collaborate and may even conflict with each other in the joint setting, naively mimicking the demonstrations will not achieve cooperation. Therefore, purely imitation learning-based approaches would not be effective. We need an approach that selectively utilizes suitable personalized expert demonstrations as guidance and allows agents to learn to cooperate via collecting reward signals from the environments.
@@
To this end, we present our algorithm, Personalized Expert-Guided MARL (PegMARL), which carries out personalized occupancy matching as a form of guidance through reward-shaping. We implement this via two discriminators. The first, a personalized behavior discriminator, evaluates local state-action pairs, providing positive incentives for actions that align with the demonstration and negative incentives for divergent ones. The second, a personalized transition discriminator, assesses whether a local state-action pair induces a desired change in dynamics similar to that observed in the demonstration, adjusting the incentive weight accordingly. We demonstrate the effectiveness of PegMARL on both discrete gridworld and continuous multi-agent particle environments (Lowe et al., 2017; Mordatch & Abbeel, 2017). The main contributions of this paper are summarized as follows:
(1) We propose PegMARL, the first approach that enables utilizing personalized demonstrations for policy learning in heterogeneous MARL environments, which (i) avoids demonstration recollection regardless of the number and type of agents involved, and (ii) is compatible with most MARL policy gradient methods.
(2) We demonstrate PegMARL’s effectiveness with personalized demonstrations in both discrete and continuous environments. Our algorithm outperforms state-of-the-art decentralized MARL algorithms, pure multi-agent imitation learning, and reward-shaping techniques in terms of scalability and convergence speed, and achieves robust performance even when provided with suboptimal demonstrations.
(3) We showcase PegMARL’s capability to also leverage joint demonstrations, regardless of whether they are sampled from co-trained or non-co-trained policies222“Co-trained policies” refers to policies that have been trained together in the same environment with shared experiences. We refer the readers to Figure 3 for an illustration and Section 6.2 of Wang et al. (2023) for more information.. Experimental results on the StarCraft environment (Samvelyan et al., 2019) demonstrate that PegMARL converges effectively even with demonstrations from non-co-trained policies, which include potentially conflicting behaviors.
@@ 2 Related Works
Imitation Learning (IL). IL methods seek to replicate an expert policy from demonstration data generated by that policy. Behavior Cloning (BC) is a commonly employed IL technique (Pomerleau, 1991; Bojarski et al., 2016), where the expert policy is estimated through supervised learning on demonstration data. However, BC is susceptible to compounding errors resulting from distribution shifts (Ross et al., 2011). Another thread of IL research is Inverse Reinforcement Learning (IRL) (Ng et al., 2000; Ziebart et al., 2008), in which the underlying reward function is estimated from the demonstration data and then used for policy learning. To alleviate the computational overhead of IRL, Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) was introduced, allowing direct policy learning from observed data without intermediate IRL steps. Song et al. (2018) extended this approach to introduce Multi-Agent GAIL (MAGAIL), which adapts GAIL to high-dimensional environments featuring multiple cooperative or competing agents. In general, IL approaches can rarely perform better than demonstrations. Therefore, they are not directly suitable for scenarios where only personalized expert demonstrations that do not demonstrate how to collaborate are available.
@@
Learning from Demonstration (LfD). In contrast to IL methods, LfD aims to leverage demonstration data to facilitate learning rather than simply mimicking expert behavior. Existing LfD works incorporate demonstration data into a replay buffer with a prioritized replay mechanism to accelerate the learning process. Due to the off-policy nature of the demonstration data, most methods are value-based (Hester et al., 2018; Vecerik et al., 2017). There has been some recent work where the demonstrations are used to aid exploration, especially in environments with large state-action spaces (Kang et al., 2018; Chen & Xu, 2022; Rengarajan et al., 2022). For instance, POfD (Kang et al., 2018) learns an implicit reward from the demonstration data using a discriminator and incorporates it into the original sparse reward. LOGO (Rengarajan et al., 2022) uses the demonstration data to directly guide the policy update: during each update iteration, the algorithm seeks a policy that closely resembles the behavior policy within a trust region. However, these approaches are primarily focused on single-agent settings.
@@
In multi-agent settings, Qiu et al. (2022) suggest using demonstrations to pretrain agents through imitation learning as a warm start, followed by optimization of the pretrained policies using standard MARL algorithms. Lee & Lee (2019) augment the experience buffer with demonstration trajectories and gradually decrease the mixing of demonstration samples during training to prevent the learned policy from being overly influenced by demonstrations. Similar to POfD, DM2 (Wang et al., 2023) enables agents to enhance their task-specific rewards by training discriminators as well. Each agent matches toward a target distribution of concurrently sampled trajectories from a joint expert policy to facilitate coordination. These approaches, however, require joint demonstrations of the same team configurations sampled from cohesively trained policies. While effective, this requirement can be cumbersome and limiting, especially in real-world scenarios where such coordinated demonstrations are difficult to obtain (as mentioned in Section 1).
Our approach differs in that we leverage only personalized expert demonstrations to learn a cooperative policy. Notably, PegMARL can also utilize joint demonstrations when they are available, including those from non-co-trained policies. By accommodating these diverse and potentially conflicting demonstrations, PegMARL can leverage a wider range of available data, offering greater flexibility and robustness across various multi-agent applications.
@@ 3 Preliminaries
We start by considering a Markov Decision Process $(\mathcal{S},\mathcal{A},P,r,\gamma)$ for a cooperative multi-agent setting with $N$ agents. Here, $\mathcal{S}$ denotes the global state across all the agents, which can be decomposed as the product of $N$ local spaces $\mathcal{S}_{i}$ as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}\times\cdots\times\mathcal{S}_{N}$. By noting the local state of agent $i$ as $s_{i}\in\mathcal{S}_{i}$, the global state is $s=(s_{1},s_{2},\cdots,s_{N})$. Similarly, we define the global action space $\mathcal{A}$ as $\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{N}$, meaning that for any $a\in\mathcal{A}$, we may write $a=(a_{1},a_{2},\cdots,a_{N})$ with $a_{i}\in\mathcal{A}_{i}$. The transition probability from state $s$ to $s^{\prime}$ after taking a joint action $a$ is denoted by $P(s^{\prime}|s,a)=\prod_{i}P_{i}(s_{i}^{\prime}|s,a)$. All agents share a common reward function $r=R(s,a)$, and $\gamma\in[0,1]$ is the discount factor. In this work, we focus on decentralized learning and define the global policy as $\pi_{\theta}(a|s)$, where $\theta\in\Theta$ are the policy parameters. Specifically, we have $\theta=(\theta_{1},\theta_{2},\cdots,\theta_{N})$ as the factorized global policy parameters, and we can write $\pi_{\theta}(a|s)=\prod_{i}\pi_{\theta_{i}}(a_{i}|s)$ using policy factorization.
@@
Personalized Tasks and MDPs. To introduce the notion of personalized demonstration, we need to extract the individual tasks of each agent from the collective tasks of multiple agents. For example, in Figure 2, the green agent’s personalized task is to open the door and the red agent’s personalized task is to reach the key without the other’s presence. We then define Personalized Markov Decision Processes (PerMDPs) $(\mathcal{S}_{i},\mathcal{A}_{i},Q_{i},r_{i},\gamma)$ for each agent or, more generally, for each $type$ of agents that share the same objective within a heterogeneous team. Here, a $type$ refers to behaviorally identical agents who share the same state and action spaces, as well as the same reward function, using the formalism from Bettini et al. (2023), and we provide the PerMDP definition for each agent for simplicity. We assume that the state space $\mathcal{S}_{i}$ and action space $\mathcal{A}_{i}$ of the personalized task $\mathcal{T}_{i}$ are the same as the local state and action spaces of agent $i$ from the joint task. The transition probability from state $s_{i}$ to $s_{i}^{\prime}$ after taking a action $a_{i}$ is represented as $Q_{i}(s^{\prime}_{i}|s_{i},a_{i})$. By following an arbitrary policy $\pi_{i}(a_{i}|s_{i})$ for this personalized task, we can collect a set of personalized demonstrations $\mathcal{B}_{i}=\{(s_{i}^{t},a_{i}^{t})\}_{t=0}^{H}$, where $H$ is the episode horizon. We’d like to emphasize that while we assume the personalized tasks and the joint task are conducted in the same environment map, the underlying transition dynamics are different. Additionally, the joint reward $r$ may not necessarily equal the summation of rewards $r_{i}$ for each personalized task.
@@ 4 MARL with Personalized Expert Demonstrations
Now, we are ready to present the main problem we are interested in solving in this work. Assume each agent $i$ is associated with a personalized task $\mathcal{T}_{i}$. We collect one set of expert demonstrations for each agent $i$ or, equivalently, each personalized task $\mathcal{T}_{i}$. By letting an expert user perform each personalized task in the respective personalized MDP, we obtain a collection of expert demonstrations denoted by $\{\mathcal{B}_{E_{1}},\mathcal{B}_{E_{2}},...,\mathcal{B}_{E_{N}}\}$. We assume that the underlying expert policy associated with $\mathcal{B}_{E_{i}}$ is $\pi_{E_{i}}(a_{i}|s_{i})$, and $\lambda^{\pi_{E_{i}}}$ is the occupancy measure following the expert’s policy $\pi_{E_{i}}$ for agent $i$. Note that while we establish the problem formulation and develop our algorithm in a fully observable setting, the experiments are conducted in both fully and partially observable settings, encompassing both discrete and continuous environments (details in Section 5).
@@ 4.1 Formulating Personalized Expert-Guided MARL
In standard Multi-Agent Reinforcement Learning, agents aim to discover optimal joint policies that maximize the long-term return $R(\pi_{\theta}):=\langle\lambda^{\pi_{\theta}},r\rangle=\frac{1}{N}\sum_{i=1}^{N}\langle\lambda^{\pi_{\theta}}_{i},r\rangle$. Typically, the learning process commences with random exploration, which is often inefficient due to MARL’s exponentially growing exploration spaces, especially when rewards are sparse. Beginning with the intuition of leveraging personalized demonstrations as guidance for how each agent should accomplish their personalized tasks to promote more effective exploration, akin to Kang et al. (2018) for single-agent cases, we can define the objective for learning from personalized demonstrations in multi-agent settings as follows:
@@
However, only imitating the personalized demonstration may not always yield favorable outcomes and can even impede the learning process. Previous works in MARL have predominantly utilized joint demonstrations as guidance. As demonstrated by DM2 (Wang et al. (2023)), trajectories must be sampled from a co-trained joint expert policy for their joint action-matching objective to converge. The crucial aspect that jointly coordinated demonstrations offer, which personalized demonstrations lack, is compatibility. To illustrate, consider the example depicted in Figure 3, where two agents aim to swap positions without entering the danger region. Two sets of optimal joint policies exist for this task: agent $a$ takes path 1 and agent $b$ takes path 2, or vice versa.
When sampling joint demonstrations from a set of co-trained policies, the agents’ behaviors will be naturally compatible. In contrast, if personalized demonstrations were provided, agents could indifferently choose between both paths to reach their goals, potentially resulting in conflicting strategies through naive imitation.
@@
Consider the scenario depicted in Figure 4, where the blue agent seeks to replicate its personalized demonstration within a multi-agent environment. The blue agent can successfully transition to the neighboring grid on the right if it’s unoccupied; otherwise, the transition fails. Viewing the blue agent as the focal agent and the green agent as part of the environment, successful imitation hinges on the agent’s “local transition” matching those of the environment where the personalized demonstrations originated. We define $P(s_{i}^{\prime}|s_{i},a_{i})$ as an approximation for the $i$-th agent’s local transition probability in the multi-agent environment, which is governed by the true transition dynamics $P(s^{\prime}|s,a)$ and the policy $\pi_{\theta}$. With this insight, we further refine the objective function as follows:
@@
where $Q(s_{i}^{\prime}|s_{i},a_{i})$ represents the true transition dynamics of the personalized MDP from which demonstrations are collected, and $\epsilon$ is a threshold parameter that theoretically controls transition mismatch tolerance. By adjusting the learning objective in this manner, the occupancy matching only happens on those local state-action pairs that guide us toward the desired next local state under the current policy.